OpenStudy (anonymous):

VERIFY (SIN^3X+COS^3X)/(SINX+COSX)=1-SINXCOSX

OpenStudy (anonymous):

Factor the numerator on the left: \[\sin^3x+\cos^3x=(\sin x+\cos x)(\sin^2x-\sin x\cos x+\cos^2x)\]

OpenStudy (anonymous):

okay.. what do I do after?

OpenStudy (anonymous):

\[\text{left}=\frac{(\sin x+\cos x)(\sin^2x-\sin x\cos x+\cos^2x)}{\sin x+\cos x}=\sin^2x-\sin x\cos x+\cos^2x\]